# Balancing Caregiving and Self-Care: Exploring Mental Health Needs of Alzheimer’s and Dementia Caregivers

### v1

@@ 5. RQ2: Practices to Address Mental Wellbeing Needs
As caregiving responsibilities and mental health challenges evolve over time, caregivers adopt a range of practices to manage their wellbeing. These practices are often ad-hoc, shaped by immediate needs and lived experiences, but can also develop into consistent routines over time. In addition, they are not static; rather, they shift in response to changing emotional, physical, and relational demands of caregiving. Section 5 provides an overview of mapping mental wellbeing practices with the different stages of caregiving. In the early stage, caregivers navigate the initial shock and uncertainty of their new responsibilities, relying on institutional resources, physical activities, and online resources to gather information and regain control. As emotional disconnection and intensified strain emerged in the middle stage, caregivers expand their coping strategies by seeking professional support, engaging in personal hobbies, participating in online support communities, setting firmer boundaries in daily life, and increasingly relying on family members for emotional and logistical assistance. In the late stage, marked by exhaustion and burnout, caregivers emphasize the importance of structured mental health interventions, stronger boundary-setting, and leveraging support from family and close friends. This temporal framing highlights the adaptive nature of wellbeing practices and underscores the need for flexible, stage-sensitive support systems that align with caregivers’ evolving emotional and practical challenges. In particular, we found two major themes of practices— 1) seeking external support and 2) adopting self-care practices.
@@
[…]
@@
[…]
@@
[…]
@@ 6.1. Need/Use of Technology
Information and Learning Resources.  Several participants mentioned how they often use a variety of technologies to get access to informational resources. They listened to podcasts (P1, P15, P22), watched YouTube videos (e.g., Dr. Natali Edmonds) (P13), and consulted sources like the Alzheimer’s Association website and Wikipedia (P5, P13). Online communities such as Reddit and Yahoo Groups provided peer advice and information on treatments (P2, P24). Some preferred short videos (P13), while others found AI tools like ChatGPT helpful for getting comprehensive, personalized answers (P20).
@@
“Even before Apple Watch, I had a Fitbit, and I counted steps on that every day and looked at it to see what I’d done. It was motivating for me. These apps are motivating for me.”—P9
@@ 6.2.1. Techno-Optimism
“It was around 3:00 AM; I was worried about what was happening with my grandfather. You can’t call the doctor at that time, so I just googled and used an AI-based tool. I explained the symptoms, and the AI advised me to keep him active. We were able to resolve the situation. The key thing about [AI] is its availability—professionals are not always available.”—P15
@@ 7. Discussion
In this study, we conducted semi-structured interviews with 25 family caregivers of individuals with AD/ADRD, and adopted inductive coding and thematic analyses (Braun and Clarke, 2019) to identify the 1) major mental health concerns of caregivers (RQ1), 2) practices and coping strategies employed by caregivers (RQ2), and 3) technologies used by caregivers in their daily caregiving and support mental wellbeing. At a high level, the interviews reinforced the motivation of our study and confirmed a critical gap in the current healthcare infrastructure and support system—caregivers’ needs are largely unaddressed, with few systematic mechanisms tailored to prioritize their mental health. In particular, AD/ADRD predominantly affects elderly individuals, whose caregiving responsibilities typically fall into one of two age groups—1) middle-aged adults caught in the “sandwich generation,” simultaneously needing to caregiving for aging parents while parenting their children, or 2) aging spouses who face their own health challenges—“I don’t want to die before my husband”—as P9 emotionally reflected about a deep-concern for their future. In this section, we discuss the implications for our work.
@@ 7.1. Theoretical Implications
In this section, we discuss how our findings are situated within existing theoretical frameworks. In particular, we contextualize our work with two relevant theoretical lenses—Social Support Behavioral Code (SSBC) (Suhr et al., 2004) and the Ethics of Care (Gilligan, 2014; Tronto, 2020).
@@ 7.1.2. Situating with the Ethics of Care
Relational Tensions and Emotional Labor.  Caregivers face significant challenges in managing family relationships and tensions, often feeling overwhelmed by caregiving responsibilities and a lack of family support. The emotional labor involved in caregiving led to additional strain. This supports prior work (Meyerhoff et al., 2022), and with the Ethics of Care (Gilligan, 2014; Tronto, 2020) emphasis on the complexities of care relationships and the moral importance of attending to these relational dynamics (Gilligan, 2014; Tronto, 2020).
@@ 7.2. Societal and Policy Implications
Our study bear implications for societal structures and policy-making, particularly in relation to financial, social, and mental health support for AD/ADRD caregivers. These insights highlight both systemic gaps and opportunities for multi-level interventions that prioritize caregiver wellbeing. We preface this discussion by acknowledging that our study was conducted in the U.S., and as such, the topics and responses reflect the socio-economic contexts of the region.
@@ 7.2.2. Combating Social Isolation Through Community-Based Interventions
Our findings reveal that caregivers experience social isolation and disrupted social lives, underscoring a need for community-based interventions. These interventions can build on and extend traditional support groups to create meaningful social connections and practical support networks. Local community centers, faith-based organizations, and neighborhood support programs could offer respite care services, organize social activities that accommodate caregivers’ scheduling constraints, and facilitate peer-to-peer connections among families facing similar challenges (Dai et al., 2023). In particular, a participant (P2) shared that their church barred their partner from attending services after the progression of AD/ADRD—significantly worsening the social lives of both of them. This reinforces for Ruitenburg et al. (2024)’s call for policy-level interventions aimed at reducing the stigma associated with dementia-related communication challenges and promoting more inclusive social environments. They suggest that current anti-stigma efforts may not sufficiently engage influential social settings such as faith-based institutions. This is a particularly critical problem to address given that the church is the most common third space in many parts, especially in rural regions, of the U.S., and being ostracized from one’s church can be akin to being ostracized from the entire local community.
@@ 7.2.3. Improving Access to Mental Health Support for Caregivers
To address these needs, accessible and proactive mental health interventions are essential, such as through counseling, peer support programs, and targeted resources for caregivers. Embedding mental health screenings into regular healthcare appointments offers a preventive approach to identifying caregivers’ distress early and intervening before crisis points (Kokorelias et al., 2019; Dayer et al., 2013). At the same time, caregivers face significant friction in accessing benefits and services due to bureaucratic complexity. Participants noted how time-consuming processes detracted their ability to focus on care. This burden could be reduced by streamlining access to public programs through simplified applications, better navigational assistance, and integrated platforms (Kim et al., 2024). This requires collaboration across healthcare providers, social services, and community organizations (Kim et al., 2024)—a system-level approach that positions caregiver wellbeing as a central, not peripheral, concern.
@@ 7.3.1. Cross-cutting Needs for Caregiver Wellbeing Technologies
Credibility.  Finally, caregivers raised concerns about the credibility of digital guidance, particularly in emotionally complex situations. P2 and P12 questioned the ability of AI tools like ChatGPT (cgp, 2024) to provide emotionally accurate or trustworthy advice. Prior work has shown that involving healthcare professionals in content development can help ensure both clinical accuracy and emotional relevance (de Ruiter et al., 2023; Sharma et al., 2024). Additionally, implementing a “source transparency” framework—where tools clearly disclose the origins of their recommendations (e.g., medical literature, professional guidelines, or clinical best practices)—can further enhance trust (Quinn et al., [n. d.]; Yoo et al., 2025b; Ehsan et al., 2023). Such transparency has been shown to improve perceived reliability in digital health systems (Kelley et al., 2024). Our findings reinforce that credibility is not just a technical issue—it is central to emotional reassurance, cognitive ease, and sustained engagement, especially as caregivers navigate high-stress and uncertain care contexts.
@@ 7.3.2. Designing for Emotionally Aware Caregiving Technologies
Building on caregivers’ evolving mental health needs, our findings point to several promising directions for emotionally supportive technologies. For instance, we can think of designing anticipatory emotional guidance tools that offer context-sensitive reflections during major care transitions—layering psychological support onto existing logistical features (Houben et al., 2024). In addition, building on identity-aware design from prior work (Ruitenburg et al., 2024), we can design self-reflective interfaces, which can help caregivers better recognize and articulate emotional strain or feelings of relational distance—as well as help with the personal growth and emotional maturity (as noticed a positive mental health effect of caregiving). We found several instances of referring to the online communities for emotional and informational support. This also support a series of prior work on the potential benefits of online support communities, both generally (Andalibi et al., 2016; Saha et al., 2020), as well as in the case of AD/ADRD caregivers (Kaliappan et al., 2025; Johnson et al., 2022). Our findings encourage design discussions on building peer support platforms that can also facilitate co-regulation of emotions, shared coping strategies, and the normalization of difficult feelings such as guilt, burnout, or grief. Finally, emotion-aware or mood-sensing interfaces can complement task-based assistance (Lee et al., 2023) by engaging with caregivers’ internal states—offering low-effort emotional check-ins or personalized nudges aligned with caregivers’ energy levels and caregiving intensity—recent research integrated behavioral sensing and generative AI capabilities for adaptive mood interventions (Nepal et al., 2024; Das Swain et al., 2025). These directions highlight the need to treat mental wellbeing not as an add-on, but as a central component of caregiver technology design.
@@ 7.3.4. Dynamic Designs: The Future of Technology to Support Caregivers’ Mental Health
Recent research in personal health informatics technology design highlights the importance of goal-aligned and adaptable systems that can support users’ shifting motivations and priorities over time (Munson et al., 2020; Sefidgar et al., 2024). Munson et al. (2020) argued that successful health tracking tools should start with an understanding of users’ personal goals and adjust as those goals evolve. Similarly, Sefidgar et al. (2024) showed how technologies that accommodate goal transitions (e.g., from symptom monitoring to emotional reflection) better align with the lived experiences of people managing chronic health conditions. Applying these insights into caregivers’ mental wellbeing, we forsee a future where mental health platforms are adaptive. Tools that incorporate stage-sensitive mental health check-ins as well as adaptable self-care prompts and interventions that respond to caregivers’ level of burnout, emotional triggers, or life transitions (for themselves and the care-recipients)—all with a goal of building good mental health skill and resilience.
@@ 7.4. Ethical Implications
As technology becomes more integrated into our lives and society, we also need to consider the ethical implications of adopting computing technologies in caregiving. Some caregivers expressed concerns about the security of personal and sensitive information. In particular, some participants desired for more personalization, however, as Pandit and Lewis described, personalization can be a double-edged sword—personalization can come at the cost of more personal data, i.e., potentially compromised privacy. This aligns with prior work on personalization-privacy tradeoffs (Asthana et al., 2024; Li et al., 2024; Zargham et al., 2022; Yoo et al., 2025a). Therefore, ensuring robust data encryption, secure storage, and compliance with regulations like the Health Insurance Portability and Accountability Act (HIPAA) is essential. Further, transparent privacy policies and giving users control over their data can enhance trust.
The recent emergence of generative AI and chatbots also brought up discussions about these AI chatbots in our interviews. In particular, the use of AI in caregiving tools raises questions about reliability and potential biases. AI systems may not fully comprehend the nuances of clinical conditions or cultural contexts, leading to inappropriate recommendations. Continuous evaluation and oversight of AI algorithms are necessary to prevent harm. It is also important to incorporate caregivers’ perspectives and needs in designing and building these chatbots—this will help utilize the values from lived experiences to anticipate AI harms and help in AI alignment (Gabriel, 2020; Shen et al., 2024; Kawakami et al., 2023). Further, there is a risk that technological solutions may widen the digital inequalities (Robinson et al., 2015). Efforts should be made to ensure that technologies are inclusive and accessible to caregivers of diverse backgrounds, including those with limited technological proficiency or resources. Finally, we highlight the ethical dilemma of relying too heavily on technology in caregiving. While technology can supplement caregiving tasks, it cannot replace the empathy and emotional understanding that human interaction offers. Therefore, striking a balance between automation and human support is crucial to maintaining the dignity and emotional wellbeing of both caregivers and care recipients.
@@ 8. Limitations and Future Directions
Our study has limitations, which also point to interesting future directions. The current study relied on a limited sample of U.S.-based caregivers and participants’ self-reports, which may not necessarily reflect the real-time AD/ADRD caregiving journey. Based on the participants’ perspective about technology for mental wellbeing, there is a clear need for future research to explore more potential uses and harms for technologies in mental wellbeing, and develop values elicitation technology that could help caregivers (Foong et al., 2024), especially in a collectivistic culture where there is an emphasis on harmonious decision-making (Ozdemir et al., 2023; Shin et al., 2013). Future research can address these gaps by incorporating a more diverse sample across cultural contexts and conducting longitudinal studies that track the use of technology for mental health over time. Expanding research to develop and evaluate personalized, adaptive digital technology-including AI-driven and emotionally sensitive support systems -could provide more tailored mental health resources (Wójcik et al., 2021) to meet the varying needs of caregivers more effectively. Future research can further explore how caregivers’ mental health needs vary across different demographic groups and caregiving contexts. For instance, while financial burden emerged as a recurring theme in our study, we did not collect explicit data on participants’ household income—a notable limitation. Examining the relationship between income levels and perceived financial strain in future work would offer valuable insights into the socioeconomic dimensions of caregiver wellbeing. Likewise, conducting participatory studies that involve other stakeholders, including community and institutional leaders, can enrich our understanding in building collaborative approaches and support strategies to cater to caregivers’ wellbeing.

### v2

@@ 5. RQ2: Practices to Address Mental Wellbeing Needs
As caregiving responsibilities and mental health challenges evolve, caregivers adopt a range of wellbeing practices. These practices are often ad-hoc, shaped by immediate needs and lived experiences, but can also develop into consistent routines over time. In addition, they are not static; rather, they shift in response to changing emotional, physical, and relational demands of caregiving. Table 4 provides an overview of mapping mental wellbeing practices with the different stages of caregiving. In the early stage, caregivers navigate the initial shock and uncertainty of their new responsibilities, relying on institutional resources, physical activities, and online resources to gather information and regain control. In the middle stage, as emotional strain deepens, they expand coping strategies—seeking professional help, joining support groups, pursuing hobbies, and leaning on family for emotional and logistical support. In the late stage, marked by burnout, caregivers emphasize structured mental health interventions, firmer boundaries, and support from close networks. This temporal framing underscores the need for flexible, stage-sensitive support systems. Across all stages, we found two major themes of practices—1) seeking external support and 2) adopting self-care practices.
@@
Physical Activities: Simple walks with the care recipient, basic stretching, deep breathing, and exercise. “We have a park nearby where we walk the dogs. We do a lot of swimming and and that sort of thing, some yoga and stretches.”—P2
Online Resources: Beginning the use of wellbeing and meditation apps, and participating in online communities to learn more about similar circumstances from community members “The [meditation] app assists me and tells me to take some deep breaths at times. So once in a while too, I also go online to edit support groups, like I mentioned earlier, so I get advice from other caregivers on managing daily challenges.”—P20
@@
Hobbies: Reading novels, playing video games. “It’s helpful because of the fact that your mind is, my mind is always, always occupied. So if I’m free, if I’m not doing anything, I’m playing video games.”—P3
Social Support from Online Communities: Regular participation in online communities, engaging with peers. “So on Reddit, there are a lot of people that caregivers, such as I. So we actually have, like, we have a group where we actually discuss so on that group, everybody comes with their own challenges and how they face them. ”—P20
Setting Boundaries in Life: Learning to decline additional responsibilities, creating physical separation, and prioritizing self-care activities. “I’ve also become better at setting my own boundaries and prioritizing my own well. Being early on, when I first started, I used to feel guilty about taking time for myself, but now I understand that some sale cases are essential for long term caregiving.”—P17
Family Support: Relying on family support to continue caregiving, and deal with difficult circumstances. “So those days I needed to, like, take care of myself, go out with friends and get things for myself, and then I also have my family come around to help, especially when it’s proven difficult”—P22
Late Stage: Emotional Exhaustion and Burnout
@@
Setting Boundaries: Major boundary decisions (facility placement), clear separation of life aspects. “Deciding to send the care recipient to memory care, more time for the caregiver.”—P19
Family support: Support from family and close friends. ”But now I’ve come to realize that I can’t do it all alone, and now I now I make a point to communicate with family or friends when I need a break, whether that having someone step in to care for a few hours ”—P21
Planning: Comprehensive care planning, professional care management. “Stay organized, prioritize tasks, take short breaks” —P17
@@ 6.1. Need/Use of Technology
Information and Learning Resources. Several participants mentioned they often used technologies to access information. They listened to podcasts (P1, P15, P22), watched YouTube videos (e.g., Dr. Natali Edmonds) (P13), and consulted sources like the Alzheimer’s Association website and Wikipedia (P5, P13). Online communities such as Reddit and Yahoo Groups provided peer advice and information on treatments (P2, P24). Some preferred short videos (P13), while others found AI tools like ChatGPT helpful for getting comprehensive, personalized answers (P20).
@@
“Even before Apple Watches, I had a Fitbit, and I counted steps on that Fitbit every day and looked at it to see what I’d done. It was motivating for me. These apps are motivating for me.”—P9
@@ 6.2.1. Techno-Optimism
“It was around 3:00 AM; I was worried about what was happening with my grandfather. You can’t call the doctor at that time, so I just googled and used an AI-based tool. I explained the symptoms and the condition, and the AI advised me to keep him active. We were able to resolve the situation. The key thing about [AI] is its availability—professionals are not always available.”—P15
@@ 7. Discussion
In this study, we conducted semi-structured interviews with 25 family caregivers of individuals with AD/ADRD, and adopted inductive coding and thematic analyses (Braun and Clarke, 2019) to identify the 1) major mental health concerns of caregivers (RQ1), 2) practices and coping strategies employed by caregivers (RQ2), and 3) technologies used by caregivers in their daily caregiving and support mental wellbeing. At a high level, the interviews reinforced the motivation of our study, and confirmed a critical gap in the current healthcare infrastructure and support system—caregivers’ needs are largely unaddressed, with few systematic mechanisms tailored to prioritize their mental health. In particular, AD/ADRD predominantly affects elderly individuals, whose caregiving responsibilities typically fall into one of two age groups—1) middle-aged adults caught in the “sandwich generation,” simultaneously needing to caregiving for aging parents while parenting their children, or 2) aging spouses who face their own health challenges—“I don’t want to die before my husband”—as P9 emotionally reflected about a deep-concern for their future. In this section, we discuss the implications for our work.
@@ 7.1. Theoretical Implications
In this section, we discuss how our findings are situated with existing theoretical frameworks. In particular, we contextualize our work with two relevant theoretical lenses—Social Support Behavioral Code (SSBC) (Suhr et al., 2004) and the Ethics of Care (Gilligan, 2014; Tronto, 2020).
@@ 7.1.2. Situating with the Ethics of Care
Relational Tensions and Emotional Labor. Caregivers face significant challenges in managing family relationships and tensions, often feeling overwhelmed by caregiving responsibilities and lack of family support. The emotional labor involved in caregiving led to additional strain. This supports prior work (Meyerhoff et al., 2022), and with the Ethics of Care (Gilligan, 2014; Tronto, 2020) emphasis on the complexities of care relationships and the moral importance of attending to these relational dynamics (Gilligan, 2014; Tronto, 2020).
@@ 7.2. Societal and Policy Implications
Our study bear implications for societal structures and policy-making, particularly in relation to financial, social, and mental health support for AD/ADRD caregivers. These insights highlight both systemic gaps and opportunities for multi-level interventions that prioritize caregiver wellbeing. We preface this discussion by acknowledging that our study was conducted in the United States and as such the topics and responses reflect the socio-economic contexts of the region.
@@ 7.2.2. Combating Social Isolation Through Community-Based Interventions
We found that caregivers often face social isolation and disrupted social lives, underscoring need for community-based interventions that go beyond traditional support groups to foster meaningful connections and practical support. Local community centers, faith-based organizations, and neighborhood support programs could offer respite care services, organize social activities that accommodate caregivers’ scheduling constraints, and facilitate peer-to-peer connections among families facing similar challenges (Dai et al., 2023). For instance, P2 shared that their church barred their partner from attending services after AD/ADRD progression—significantly worsening the social lives of both of them. This echoes for Ruitenburg et al. (2024)’s call for policy-level interventions aimed at reducing the stigma associated with dementia-related stigma and inclusion in key social spaces. As faith-based institutions and churches often serve as vital community hubs—especially in rural areas—exclusion can lead to broader social ostracism.
@@ 7.2.3. Improving Access to Mental Health Support for Caregivers
To address these needs, accessible and proactive mental health interventions are essential, such as counseling, peer support programs, and targeted resources for caregivers. Embedding mental health screenings into regular healthcare visits can help detect distresses early and prevent crises (Kokorelias et al., 2022; Dayer et al., 2013). At the same time, caregivers face significant friction in accessing benefits and services due to bureaucratic complexity. Participants noted how time-consuming processes detract from caregiving. Streamlining access to public programs through simplified applications, better guidance, and integrated platforms can ease this burden (Kim et al., 2024). This requires collaboration across healthcare providers, social services, and community organizations (Kim et al., 2024)—a system-level approach that positions caregiver wellbeing as a central, not peripheral, concern.
@@ 7.3.1. Cross-cutting Needs for Caregiver Wellbeing Technologies
Credibility. Finally, caregivers raised concerns about the credibility of digital guidance, particularly in emotionally complex situations. P2 and P12 questioned the ability of AI tools like ChatGPT (cgp, 2024) to provide emotionally accurate or trustworthy advice. Prior work has shown that involving healthcare professionals in content development can help ensure both clinical accuracy and emotional relevance (de Ruiter et al., 2023; Sharma et al., 2024). Additionally, implementing a “source transparency” framework—where tools clearly disclose the origins of their recommendations (e.g., medical literature, professional guidelines, or clinical best practices)—can further enhance trust (Ehsan et al., 2023; Quinn et al., [n. d.]; Yoo et al., 2025b). Such transparency has been shown to improve perceived reliability in digital health systems (Kelley et al., 2024). Our findings reinforce that credibility is not just a technical issue—it is central to emotional reassurance, cognitive ease, and sustained engagement, especially as caregivers navigate high-stress and uncertain care contexts.
@@ 7.3.2. Designing for Emotionally Aware Caregiving Technologies
Building on caregivers’ evolving mental health needs, our findings point to several promising directions for emotionally supportive technologies. For instance, we can think of designing anticipatory emotional guidance tools that offer context-sensitive reflections during major care transitions—layering psychological support onto existing logistical features (Houben et al., 2024). In addition, building on identity-aware design from prior work (Ruitenburg et al., 2024), we can design self-reflective interfaces, which can help caregivers better recognize and articulate emotional strain or feelings of relational distance—as well as help with the personal growth and emotional maturity (as noticed a positive mental health effect of caregiving). We found several instances of referring to the online communities for emotional and informational support. This also support a series of prior work on the potential benefits of online support communities, both generally (Andalibi et al., 2016; Saha et al., 2020), as well as in the case of AD/ADRD caregivers (Kaliappan et al., 2025; Johnson et al., 2022). Our findings encourage design discussions on building peer support platforms that can also facilitate co-regulation of emotions, shared coping strategies, and the normalization of difficult feelings such as guilt, burnout, or grief. Finally, emotion-aware or mood-sensing interfaces can complement task-based assistance (Lee et al., 2023) by engaging with caregivers’ internal states—offering low-effort emotional check-ins or personalized nudges aligned with caregivers’ energy levels and caregiving intensity—recent research integrated behavioral sensing and generative AI capabilities for adaptive mood interventions (Das Swain et al., 2025; Nepal et al., 2024). These directions highlight the need to treat mental wellbeing not as an add-on, but as a central component of caregiver technology design.
@@ 7.3.4. Dynamic Designs: The Future of Technology to Support Caregivers’ Mental Health
Recent research in personal health informatics technology design highlights the importance of goal-aligned and adaptable systems that can support users’ shifting motivations and priorities over time (Munson et al., 2020; Sefidgar et al., 2024). Munson et al. (2020) argued that successful health tracking tools should start with an understanding of users’ personal goals and adjust as those goals evolve. Similarly, Sefidgar et al. (2024) showed how technologies that accommodate goal transitions (e.g., from symptom monitoring to emotional reflection) better align with the lived experiences of people managing chronic health conditions. Applying these insights into caregivers’ mental wellbeing, we foresee a future where mental health platforms are adaptive. Tools that incorporate stage-sensitive mental health check-ins as well as adaptable self-care prompts and interventions that respond to caregivers’ level of burnout, emotional triggers, or life transitions (for themselves and the care-recipients)—all with a goal of building good mental health skills and resilience.
@@ 7.4. Ethical Implications
As technology becomes more integrated into our lives and society, we also need to consider the ethical implications of adopting computing technologies in caregiving. Some caregivers expressed concerns about the security of personal and sensitive information. In particular, some participants desired for more personalization, however, as Pandit and Lewis described, personalization can be a double-edged sword—personalization can come at the cost of more personal data, i.e., potentially compromised privacy. This aligns with prior work on personalization-privacy tradeoffs (Asthana et al., 2024; Li et al., 2024; Zargham et al., 2022). Therefore, ensuring robust data encryption, secure storage, and compliance with regulations like the Health Insurance Portability and Accountability Act (HIPAA) is essential. Further, transparent privacy policies and giving users control over their data can enhance trust.
The recent emergence of generative AI and chatbots also brought up discussions about these AI chatbots in our interviews. In particular, the use of AI in caregiving tools raises questions about reliability and potential biases. AI systems may not fully comprehend the nuances of clinical conditions or cultural contexts, leading to inappropriate recommendations. Continuous evaluation and oversight of AI algorithms are necessary to prevent harm. It is also important to incorporate caregivers’ perspectives and needs in designing and building these chatbots—this will help utilize the values from lived experiences to anticipate AI harms and help in AI alignment (Gabriel, 2020; Shen et al., 2024; Saha et al., 2025; Kawakami et al., 2023; Shi et al., 2025). Further, there is a risk that technological solutions may widen the digital inequalities (Robinson et al., 2015). Efforts should be made to ensure that technologies are inclusive and accessible to caregivers of diverse backgrounds, including those with limited technological proficiency or resources. Finally, we highlight the ethical dilemma of relying too heavily on technology in caregiving. While technology can supplement caregiving tasks, it cannot replace the empathy and emotional understanding that human interaction offers. Therefore, striking a balance between automation and human support is crucial to maintaining the dignity and emotional wellbeing of both caregivers and care recipients.
@@ 8. Limitations and Future Directions
Our study has limitations, which also point to interesting future directions. The current study relied on a limited sample of U.S.-based caregivers and relied on participants’ self-reported, which may not necessarily reflect the real-time AD/ADRD caregiving journey. Based on the participants’ perspective about technology for mental wellbeing, there is a clear need for future research to explore more potential uses and harms for technologies in mental wellbeing, and develop values elicitation technology that could help caregivers (Foong et al., 2024), especially in a collectivistic culture where there is an emphasis on harmonious decision-making (Ozdemir et al., 2023; Shin et al., 2013). Future research can address these gaps by incorporating a more diverse sample across cultural contexts and conducting longitudinal studies that track the use of technology for mental health over time. Expanding research to develop and evaluate personalized, adaptive digital technology-including AI-driven and emotionally sensitive support systems -could provide more tailored mental health resources (Wójcik et al., 2021) to meet the varying needs of caregivers more effectively. Future research could further explore how caregivers’ mental health needs vary across different demographic groups and caregiving contexts. For instance, while financial burden emerged as a recurring theme in our study, we did not collect explicit data on participants’ household income—a notable limitation. Examining the relationship between income levels and perceived financial strain in future work would offer valuable insights into the socioeconomic dimensions of caregiver wellbeing. Likewise, conducting participatory studies that involve other stakeholders, including community and institutional leaders, can enrich our understanding in building collaborative approaches and support strategies to cater to caregivers’ mental wellbeing.